# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Periodically driven Floquet quantum many-body systems have revealed new insights into the rich interplay of thermalization, and growth of entanglement. The phenomenology of dynamical freezing, whereby a translationally invariant many-body system exhibits emergent conservation laws and a slow growth of entanglement entropy at certain fixed ratios of a drive amplitude and frequency, presents a novel paradigm for retaining memory of an initial state up to late times. Previous studies of dynamical freezing have largely been restricted to a high-frequency Floquet-Magnus expansion, and numerical exact diagonalization, which are unable to capture the slow approach to thermalization (or lack thereof) in a systematic fashion. By employing Floquet flow-renormalization, where the time-dependent part of the Hamiltonian is gradually decoupled from the effective Hamiltonian using a sequence of unitary transformations, we unveil the universal approach to dynamical freezing and beyond, at asymptotically late times. We analyze the fixed-point behavior associated with the flow-renormalization at and near freezing using both exact-diagonalization and tensor-network based methods, and contrast the results with conventional prethermal phenomenon. For a generic non-integrable spin Hamiltonian with a periodic cosine wave drive, the flow approaches an unstable fixed point with an approximate emergent symmetry. We observe that at freezing the thermalization timescales are delayed compared to away from freezing, and the flow trajectory undergoes a series of instanton events. Our numerical results are supported by analytical solutions to the flow equations.
@@ I Introduction
Figure 1: A schematic depiction of the Floquet flow-renormalization trajectory (blue) along the fRG time, $\lambda$. The insets show the evolution of many-body energy-levels between the different fixed points, e.g. from $\lambda=0$ to the first (unstable) prethermal fixed-point (‘$P$’), followed by a sequence of instanton events that connect intermediate fixed points ($I_{1},~{}I_{2},...$), and final approach to the thermal (‘$T$’) fixed-point. The flow starts with the bare Hamiltonian (many-body bandwidth, $W$), and a Floquet drive with frequency, $\Omega$. The flow to $P$ leads to a renormalization of the many-body spectrum. The energy-levels associated with the effective Hamiltonian at the thermal fixed-point $T$ are folded into an interval shorter than $\Omega$.
@@
In parallel, a number of studies based primarily on numerical exact diagonalization (ED) and supporting Floquet-Magnus expansion, have reported the emergence of approximate conservation laws and slow dynamics in generic translationally invariant, periodically driven spin models at special ratios of the drive amplitude and frequency. This distinct route to avoiding Floquet thermalization has been denoted “Dynamical Freezing” [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Recent works by some of us have extended this phenomenology to a strongly interacting fermionic setup analyzed using Keldysh field-theoretic methods [54], and interacting bosonic models with a large local Hilbert-space dimension where a semiclassical correspondence reveals suppression of the chaotic Lyapunov exponent at the frozen points  [55]. Unfortunately, a vast majority of analytical attempts to study dynamical freezing have relied on the uncontrolled and perturbative Magnus expansion in a rotating frame, which does not produce a convergent series for generic interacting systems, and potentially misses important non-perturbative effects relevant for addressing fundamental questions related to thermalization (or lack thereof) at freezing. Additionally, the numerical studies have largely been restricted to ED on relatively modest system-sizes upto finite times, with the exception of one recent study for a very specific (i.e. square-wave) drive protocol [53]. This has left a number of basic questions about dynamical freezing for more generic drives unanswered, which will be addressed definitively in this work.
@@
First, it is natural to question what determines the location of the dynamically frozen points. For spin models with low-dimensional Hilbert spaces, the location typically correlates with the values of ratio of drive amplitude and frequency where the leading-order Magnus correction vanishes. However, some of us have previously demonstrated a quantitative shift in these locations due to the higher-order Magnus corrections in models with a larger on-site Hilbert-space dimension [55]. Finding an alternative method to determine the freezing points, without just relying on a combination of numerics and the Magnus expansion is a desirable goal. Second, and more importantly, whether dynamical freezing is exact and robust in the thermodynamic system-size and late-time limit, and if not, what controls the possible decay away from this special behavior has not been determined beyond the same limitations imposed by exact-diagonalization and Magnus-expansion for generic drives.
@@
In this work, we improve on these limitations of prior studies and provide a universal description of dynamical freezing based on a Floquet flow-renormalization group (fRG) framework  [56, 57, 58, 59, 60, 61, 62, 63, 64]. Fundamentally, the formalism generates a sequence of infinitesimal unitary transformations to systematically eliminate the time-dependent part of the driven Hamiltonian, resulting in a flow equation for the different coupling constants in the effective model. The method in the equilibrium setting is a generalization of the more widely known Schrieffer-Wolff transformations [65]; see Appendix A for a brief review of this connection. As we will demonstrate in this paper by focusing on a generic, non-integrable one-dimensional spin model with a “cosine-wave” drive, the flow trajectory along the fRG time at the putative freezing-point organizes into a sequence of unstable fixed-points, except the last, thermal (‘$T$’) fixed-point; see Fig. 1 for a schematic depiction. Along the flow trajectory, the approach to the first fixed-point starting from the bare Hamiltonian is of special interest. It is an unstable prethermal (‘$P$’) fixed-point, associated with the emergence of an approximately conserved effective Hamiltonian, valid up to an exponentially long timescale in the drive frequency, $\Omega$. Subsequently, the systems flows to the thermal fixed point via a sequence of intermediate fixed points $I_{1},I_{2},\dots$. The flow in between these intermediate fixed points show characteristics of instanton events.
@@
We will demonstrate explicitly that dynamical freezing manifests as an additional emergent conservation law associated with the prethermal fixed point. Exact dynamical freezing would correspond to a vanishing operator norm for the commutator between the emergent conserved quantity and the undriven part of the Hamiltonian near the prethermal fixed point. Generically, we find that the above commutator norm does not vanish exactly at freezing, but exhibits a non-trivial behavior as a function of the drive frequency, system size, and fRG time. Specifically, the emergent symmetry becomes more pronounced as the high frequency limit $\Omega\to\infty$ is approached. From that perspective, dynamical freezing is asymptotic in the high-frequency limit. Additionally, given that in a generic non-integrable system, the prethermal fixed point is expected to be unstable towards eventual thermalization, it is not surprising that the flow will eventually divert to the thermal fixed point.
@@
In this work, we have also analyzed carefully the flow from the unstable prethermal fixed point to the thermal fixed point by combining powerful numerical and analytical techniques. Interestingly, we find that the system flows via a sequence of “fixed-points”, where the intermediate evolution in fRG time shows characteristics of instanton-like events. The renormalized effective Hamiltonian in the vicinity of these intermediate fixed-points is approximately constant and the associated flow is exponentially slow in $\exp(-\Omega/J)$, where $J$ is the typical single-particle energy scale of the system. The transition between two fixed points occurs rapidly in a “short” window of fRG time, which is on the order of $J^{-1}$. We associate these “tunneling” type events as instantons (see Fig. 1), which are fundamentally non-perturbative and lie beyond the reach of more conventional perturbative methods, which includes the Floquet-Dyson perturbation theory [45] or the Floquet-Magnus expansion [66], but are readily accessible via the flow formalism. These perturbative methods are expected to describe the flow from the bare Hamiltonian to the prethermal fixed point. In contrast, the instantons signal a departure from the prethermal fixed point and the breakdown of perturbative expansions, which ultimately captures the universal flow to the thermal fixed-point.
@@
The rest of the paper is organized as follows: In Sec. II, we review the Floquet-flow formalism and the flow equation. We also introduce the characteristic features associated with dynamical freezing, as manifest in the flow formalism. In Sec. III, we use a simple example of a driven harmonic oscillator to illustrate the flow towards putative freezing using the flow formalism. In Sec. IV, we turn to the solutions of the flow equations for a non-integrable spin chain driven by a cosine-wave first using both exact diagonalization and matrix-product operator based methods. In Sec. V, we turn to a general analytical treatment of the flow equation that applies to models beyond the specific spin chain Hamiltonian, which can describe the early-time approach to the unstable prethermal fixed-point, as well as the late-time thermalization due to instantons. We end with an outlook in Sec. VI. The appendices contain a number of technical details on the methods used in this paper.
@@ II Floquet-flow renormalization and
We review the flow formalism, developed originally by Wegner, Wilson and others [67, 68] and adapted to a variety of strongly correlated problems [56] (including problems beyond the equilibrium setting [58, 59, 60, 57, 62, 63, 64, 69, 70]). The discussion here closely resembles the one introduced in Ref. [63]. We are interested in time-dependent Hamiltonians expressed as,
@@
where $H_{-1}=H_{1}^{\dagger}$. Treating Eq. (1) as the initial state of a renormalization process, the flow generates a sequence of infinitesimal unitary transformations, preserving the spectrum of the stroboscopic unitary evolution operator (see Appendix B for numerical details) [71],
@@
where $\lambda$ represents the “fRG time” along the flow trajectory. Note that $\lambda$ enters here dimensionally as real time with dimension $\sim$[time] in contrast to the conventional Wegner-Wilson picture [67, 68], where $\lambda\sim$ [time]2.
We solve the coupled matrix equations, Eq. (3a) and Eq. (3b), using two approaches. First, for small system sizes we employ exact-diagonalization by representing the operators explicitly in the many-body Hilbert space and solving the equations exactly using the fourth-order Runge-Kutta (RK4) method [57]. The second approach is tailored for larger system sizes and utilizes matrix product operators (MPOs) [63]. MPOs encode general many-body operators as products of operator-valued matrices, with each matrix acting locally at a single site. Once the operators $H_{0}(\lambda)$ and $H_{1}(\lambda)$ are expressed in the MPO form, we again solve the equations using the Runge-Kutta method. MPOs facilitate operations such as addition, evaluation, and computation of commutators, yielding exponentially localized MPOs [72, 73]. MPOs have been applied to solve flow equations for many-body localized Hamiltonians [60, 74, 75]. All ED results are obtained using the QuSpin [76, 77] Python package, while the MPO-based computations are performed using the Julia package ITensor [78, 79, 80].
@@
$t_{\text{eff}}(\lambda)\sim\left\|H_{1}(\lambda)\right\|^{-1}{\left\|{\hat{1}}\right\|}.$ (5)
Here ${\left\|{M}\right\|}^{2}={\rm Tr}(M^{\dagger}M)$ denotes the Frobenius norm of the operator, and Eq. (5) is normalized by the norm of the identity operator $\hat{1}$ on the many-body Hilbert space.
The typical behavior of the flow equation Eq. (3) can be divided into three regimes:
@@
    For fRG time, $0\leq\lambda\lesssim\Omega^{-1}$, the flow is in a ramp-up regime where $H_{0}(\lambda)$ is rapidly renormalized by the drive. In the high-frequency limit, this ramp-up flow can be described perturbatively using an expansion similar to the Floquet-Magnus expansion (see Sec. V and Appendix C).
@@
    Dynamical freezing manifests as an approximate emergent symmetry of $H_{0}(\lambda)$ in the prethermal regime, for all $\Omega^{-1}\lesssim\lambda\lesssim\lambda_{\text{min}}$. Consider the emergent symmetry associated with the bare drive component in the Hamiltonian, $H_{1}(0)=H_{-1}(0)=A{\cal{Q}}$, where ${\cal{Q}}$ is the “charge” associated with the emergent symmetry, and $A$ is the drive amplitude. The freezing phenomenon occurs in the fast and strong drive limit (i.e. large $A,\Omega$), where the ratio $A/\Omega$ takes discrete special values. At the freezing point, the prethermal Hamiltonian $H_{0}(\lambda)$ shows an emergent approximate symmetry, characterized by the commutator $[H_{0}(\lambda),{\cal{Q}}]\sim\mathcal{O}(1/\Omega^{2})$, which vanishes in the $\Omega\to\infty$ limit. Therefore, freezing is only approximate at any finite $\Omega$, and importantly, does not survive when the system flows out of the prethermal regime.
@@
    When $\lambda\gtrsim\lambda_{\text{min}}$, the system enters the thermalization regime, where $H_{0}(\lambda)$ eventually flows to a “random” matrix, signifying the heating to infinite temperature. From the explicit integration of the flow equations, we find that this complex process occurs through a series of instanton events, where $H_{1}(\lambda)$ becomes sizable in a narrow window of fRG time, and in the same window a pair of eigenvalues of $H_{0}(\lambda)$ approach each other. As a result of the instanton events, eigenvalues of $H_{0}(\lambda)$ will ultimately be pushed into an interval of width $\Omega$, which means $H_{0}(\lambda\to\infty)$ is now interpreted as a Floquet Hamiltonian. A more detailed technical discussion of these timescales are presented in Sec. V and Appendix D.

### v2

# Floquet-Thermalization via Instantons near Dynamical Freezing

@@ Abstract
Periodically driven Floquet quantum many-body systems have revealed new insights into the rich interplay of thermalization, and growth of entanglement. The phenomenology of dynamical freezing, whereby a translationally invariant many-body system exhibits emergent conservation laws and a slow growth of entanglement entropy at certain fixed ratios of a drive amplitude and frequency, presents a novel paradigm for retaining memory of an initial state up to late times. Previous studies of dynamical freezing have largely been restricted to a high-frequency Floquet-Magnus expansion, and numerical exact diagonalization. Both techniques are unable to capture the slow approach to thermalization, or lack thereof, in a systematic fashion. By employing Floquet flow-renormalization, where the time-dependent part of the Hamiltonian is gradually decoupled from the effective Hamiltonian using a sequence of unitary transformations, we unveil the universal approach to dynamical freezing and beyond, at asymptotically late times. We analyze the fixed-point behavior associated with the flow-renormalization at and near freezing using both exact-diagonalization and tensor-network based methods, and contrast the results with conventional prethermal phenomenon. For a generic non-integrable spin Hamiltonian with a periodic cosine wave drive, the flow approaches an unstable fixed point with an approximate emergent symmetry. We observe that at freezing the thermalization timescales are delayed compared to away from freezing, and the flow trajectory undergoes a series of instanton events. Our numerical results are supported by analytical solutions to the flow equations.
@@ I Introduction
Figure 1: A schematic depiction of the Floquet flow-renormalization trajectory (blue) along the fRG time, $\lambda$. The insets show the evolution of many-body energy-levels between the different fixed points, e.g. from $\lambda=0$ to the first (unstable) prethermal fixed-point (‘$P$’), followed by a sequence of instanton events that connect intermediate fixed points ($I_{1},~I_{2},$...), and final approach to a possibly thermal (‘$T$’) fixed-point, depending on a non-trivial interplay of the drive frequency and finite system size. The flow starts with the bare Hamiltonian (many-body bandwidth, $W$), and a Floquet drive with frequency, $\Omega$. The flow to $P$ leads to a renormalization of the many-body spectrum. The energy-levels associated with the effective Hamiltonian at the thermal fixed-point $T$ are folded into an interval smaller than $\Omega$. See Table 1 for a relationship between the fRG time, $\lambda$, and real time, $t$, along with a summary of the different regimes encountered during the flow vis-à-vis these timescales.
@@
In parallel, a number of studies based primarily on numerical exact diagonalization (ED) and supporting Floquet-Magnus expansion, have reported the emergence of approximate conservation laws and slow dynamics in generic translationally invariant, periodically driven spin models at special ratios of the drive amplitude and frequency. This distinct route to avoiding Floquet thermalization has been denoted “Dynamical Freezing” [21, 11, 39, 63, 37, 49, 38, 70, 5, 29, 7, 30, 8, 64, 36]. Recent works by some of us have extended this phenomenology to a strongly interacting fermionic setup analyzed using Keldysh field-theoretic methods [33], and interacting bosonic models with a large local Hilbert-space dimension where a semiclassical correspondence reveals suppression of the chaotic Lyapunov exponent at the frozen points  [50]. Unfortunately, a vast majority of analytical attempts to study dynamical freezing have relied on the uncontrolled and perturbative Floquet-Magnus expansion in a rotating frame, which does not produce a convergent series for generic interacting systems, and potentially misses important non-perturbative effects relevant for addressing fundamental questions related to thermalization (or lack thereof) at freezing. This has left a number of basic questions about dynamical freezing for more generic drives unanswered, which will be addressed definitively in this work. Additionally, the numerical studies have largely been restricted to ED on relatively modest system-sizes upto finite times. The only notable exception is the recent numerical study for an infinite system size, but carried out for only a square-wave drive protocol [36]. In contrast, the focus of this manuscript will be on the fate of a Floquet many-body system with a sinusoidal drive — which is known to exhibit distinct behavior even within the perturbative Floquet-Magnus expansion [50] — but now obtained using a number of complementary numerical and analytical non-pertubative methods.
@@
First, it is natural to question what determines the location of the dynamically frozen points. For spin models with low-dimensional Hilbert spaces, the location typically correlates with the values of ratio of drive amplitude and frequency where the leading-order Magnus correction vanishes. However, some of us have previously demonstrated a quantitative shift in these locations due to the higher-order Magnus corrections in models with a larger on-site Hilbert-space dimension [50]. Finding an alternative method to determine the freezing points, without relying on a combination of numerics and the Floquet-Magnus expansion is a desirable goal. Second, and more importantly, whether dynamical freezing is exact and robust in the thermodynamic system-size and late-time limit, and if not, what controls the possible decay away from this special behavior has not been determined beyond the same limitations imposed by exact-diagonalization and Magnus-expansion for generic drives.
@@
In this work, we improve on these limitations of prior studies and provide a universal description of dynamical freezing based on a Floquet flow-renormalization group (fRG) framework  [41, 81, 55, 60, 91, 88, 78, 18, 77]. Fundamentally, the formalism generates a sequence of infinitesimal unitary transformations to systematically eliminate the time-dependent part of the driven Hamiltonian, resulting in a flow equation for the different coupling constants in the effective model. The method in the equilibrium setting is a generalization of the more widely known Schrieffer-Wolff transformations [14]; see Appendix A for a brief review of this connection and Appendix B for details on its numerical implementation. As we will demonstrate in this paper by focusing on a generic, non-integrable one-dimensional spin model with a “cosine-wave” drive, the flow trajectory along the fRG time at the putative freezing-point organizes into a sequence of unstable fixed-points, except the possibly inevitable thermal (‘$T$’) fixed-point that appears at asymptotically late times  [^1] For finite systems, this final fixed point appears “thermal” only when the drive frequency $\Omega<\Omega_{*}(L)$, where $\Omega_{*}(L)$ is a system-size ($L$) dependent threshold frequency; see Sec. II and V for more details. ; see Fig. 1 for a schematic depiction. Along the flow trajectory, the approach to the first fixed-point starting from the bare Hamiltonian is of special interest. It is an unstable prethermal (‘$P$’) fixed-point, associated with the emergence of an approximately conserved effective Hamiltonian, valid up to an exponentially long timescale in the drive frequency, $\Omega$. Subsequently, the system flows via a complex sequence of intermediate fixed points $I_{1},I_{2},\dots$. The flow in between these intermediate fixed points show characteristics of instanton events. However, as we will discuss below, it is important to note at the outset that capturing the flow to the thermal fixed-point numerically is exceptionally challenging given the exponential timescales involved, and remains an exciting arena for future research.
@@
We will demonstrate explicitly that dynamical freezing manifests as an additional emergent conservation law associated with the prethermal fixed point. Exact dynamical freezing would correspond to a vanishing operator norm for the commutator between the emergent conserved quantity and the undriven part of the Hamiltonian near the prethermal fixed point. Generically, we find that the above commutator norm does not vanish exactly at freezing, but exhibits a non-trivial behavior as a function of the drive frequency, system size, and fRG time. Specifically, the emergent symmetry becomes more pronounced as the high frequency limit $\Omega\to\infty$ is approached. From that perspective, dynamical freezing is asymptotic in the high-frequency limit.
@@
In this work, we have also analyzed carefully the flow awayfrom the unstable prethermal fixed point by combining powerful numerical and analytical techniques. Interestingly, we find that the system flows via a sequence of “fixed-points”, where the intermediate evolution in fRG time shows characteristics of instanton-like events. The renormalized effective Hamiltonian in the vicinity of these intermediate fixed-points is approximately constant and the associated flow is exponentially slow in $\exp(-\Omega/J)$, where $J$ is the typical single-particle energy scale of the system. The transition between two fixed points occurs rapidly in a “short” window of fRG time, which is on the order of $J^{-1}$. We associate these “tunneling” type events as instantons (see Fig. 1), which are fundamentally non-perturbative and lie beyond the reach of more conventional perturbative methods, which includes the Floquet-Dyson perturbation theory [38] or the Floquet-Magnus expansion [24], but are readily accessible via the flow formalism. These perturbative methods are expected to describe the flow from the bare Hamiltonian to the prethermal fixed point. In contrast, the instantons signal a departure from the prethermal fixed point and the breakdown of perturbative expansions, which ultimately must capture the universal flow to the thermal fixed-point. We note that capturing and diagnosing the full flow towards this “latest” time thermalization for any appreciable system size is an exceedingly slow process, given the exponentially long timescales involved, and is therefore well beyond the scope of any numerical method. However, as we discuss below, our results associated with the sequence of instanton-like events suggest that the ultimate fate and fixed-point behavior is tied to this slow thermalization for drive frequencies that lie below a system-size dependent threshold.
@@
The rest of the paper is organized as follows: In Sec. II, we review the Floquet-flow formalism and the flow equation. We also introduce the characteristic features associated with dynamical freezing, as manifest in the flow formalism. In Sec. III, we use a simple example of a driven harmonic oscillator to illustrate the flow towards putative freezing using the flow formalism. In Sec. IV, we turn to the solutions of the flow equations for a non-integrable spin chain driven by a cosine-wave first using both exact diagonalization and matrix-product operator based methods. In Sec. V, we present results for a number of numerical diagnostics that characterize the approach to thermalization near dynamical freezing as a function of system size and drive frequency. In Sec. VI, we turn to a general analytical treatment of the flow equation that applies to models beyond the specific spin chain Hamiltonian, which can describe the early-time approach to the unstable prethermal fixed-point, as well as the late-time thermalization due to instantons. We end with an outlook in Sec. VII. The appendices contain a number of technical details on the methods used in this paper. In App. A, we review the connection between Schrieffer-Wolff transformations and the flow formalism. In App. B, we describe some additional numerical details for simulating the flow equations. In App. C, we present some additional numerical results supporting conclusions in the main text. In App. D, we present details of analytical analysis of the flow equation in the early time regime, which behaves similarly as the Floquet-Magnus expansion. In App. E, we estimate the time scales associated with thermalization in the flow formalism. In App. F, we review the Floquet-Magnus expansion of the spin chain Hamiltonian we used.
@@ II Floquet-flow renormalization and
We review the flow formalism, developed originally by Wegner, Wilson and others [82, 31] and adapted to a variety of strongly correlated problems [41] (including problems beyond the equilibrium setting [55, 60, 91, 81, 78, 18, 77, 68, 79, 15]). The discussion here closely resembles the one introduced in Ref. [18], which is reviewed in greater detail in Appendix A. We are interested in time-dependent Hamiltonians expressed as,
@@
where $H_{-1}=H_{1}^{\dagger}$. Treating Eq. (1) as the initial state of a renormalization process, the flow generates a sequence of infinitesimal unitary transformations, preserving the spectrum of the stroboscopic unitary evolution operator [16],
@@
where $\lambda$ represents the “fRG time” along the flow trajectory. Note that $\lambda$ enters here dimensionally as real time with dimension $\sim$[time] in contrast to the conventional Wegner-Wilson picture [82, 31], where $\lambda\sim$ [time]2. See Table 1 for a glossary of the different timescales that describe the dynamics within the flow RG framework.
We solve the coupled matrix equations, Eq. (3a) and Eq. (3b), using two approaches. First, for small system sizes we employ exact-diagonalization by representing the operators explicitly in the many-body Hilbert space and solving the equations exactly using the fourth-order Runge-Kutta (RK4) method [81]. In the supplement [1], we provide numerical evidences that Eq. (3) does preserve the eigenvalues of the Floquet Hamiltonian. The second approach is tailored for larger system sizes and utilizes matrix product operators (MPOs) [18]. MPOs encode general many-body operators as products of operator-valued matrices, with each matrix acting locally at a single site. Once the operators $H_{0}(\lambda)$ and $H_{1}(\lambda)$ are expressed in the MPO form, we again solve the equations using the Runge-Kutta method. MPOs facilitate operations such as addition, evaluation, and computation of commutators, yielding exponentially localized MPOs [13, 89]. MPOs have been applied to solve flow equations for many-body localized Hamiltonians [91, 80, 76]. All ED results are obtained using the QuSpin [85, 84] Python package, while the MPO-based computations are performed using the Julia package ITensor [10, 26, 25]. See Appendix B for additional details.
@@
$t_{\text{eff}}(\lambda)\sim\left\|H_{1}(\lambda)\right\|^{-1}{\left\|{\hat{1}}\right\|}$, (5)
which will provide useful estimates on the (pre-)thermalization timescales for our driven problem, as we will describe below. Here ${\left\|{M}\right\|}^{2}={\rm Tr}(M^{\dagger}M)$ denotes the Frobenius norm of the operator, and Eq. (5) is normalized by the norm of the identity operator $\hat{1}$ on the many-body Hilbert space.
The typical behavior of the flow equation Eq. (3) can be divided into three regimes (see also Table 1):
@@
    For fRG time, $0\leq\lambda\lesssim\Omega^{-1}$, the flow is in a ramp-up regime where $H_{0}(\lambda)$ is rapidly renormalized by the drive. In the high-frequency limit, this ramp-up flow can be described perturbatively using an expansion similar to the Floquet-Magnus expansion (see Sec. VI).
@@
    Dynamical freezing manifests as an approximate emergent symmetry of $H_{0}(\lambda)$ in the prethermal regime, for all $\Omega^{-1}\lesssim\lambda\lesssim\lambda_{\text{min}}$. Consider the emergent symmetry associated with the bare drive component in the Hamiltonian, $H_{1}(0)=H_{-1}(0)=A{\cal{Q}}$, where $\cal{Q}$} is the “charge” associated with the emergent symmetry, and $A$ is the drive amplitude. The freezing phenomenon occurs in the fast and strong drive limit (i.e. large $A,\Omega$), where the ratio $A/\Omega$ takes discrete special values. At the freezing point, the prethermal Hamiltonian $H_{0}(\lambda)$ shows an emergent approximate symmetry, characterized by the commutator $H_{0}(\lambda),{\cal{Q}}]\sim\mathcal{O}(1/\Omega^{2})$, which vanishes in the $\Omega\to\infty$ limit. Therefore, freezing is only approximate at any finite $\Omega$, and importantly, does not survive when the system flows out of the prethermal regime.
@@
    When $\lambda\gtrsim\lambda_{\text{min}}$, the system enters the thermalization regime, where $H_{0}(\lambda)$ flows to a more non-local Hamiltonian. When the drive frequency $\Omega$ is smaller than a system-size dependent threshold $\Omega_{*}(L)$, the system eventually flows to a “thermal” fixed point, where the operator-entanglement entropy is close to that of a random-matrix [93]. Based on our numerical results on small system sizes (see Sec. V), we find that $\Omega_{*}(L)$ weakly increases with $L$. Whether the threshold diverges in the thermodynamic limit or saturates to a finite value will determine the eventual fate of freezing, which remains an exciting direction to analyze further via new numerical developments in future work. From the explicit integration of the flow equations, we find that this complex process occurs through a series of instanton events, where $H_{1}(\lambda)$ becomes sizable in a narrow window of fRG time, and in the same window a pair of eigenvalues of $H_{0}(\lambda)$ approach each other. As a result of the instanton events, eigenvalues of $H_{0}(\lambda)$ will ultimately be pushed into an interval of width $\Omega$, which means $H_{0}(\lambda\to\infty)$ is now interpreted as a Floquet Hamiltonian. A more detailed technical discussion of these timescales are presented in Sec. VI and Appendix E.
@@
fRG time ($\lambda$) Physical Significance Real time ($t$) Physical Significance
Ramp-up regime: $\mathcal{O}(1)\times\Omega^{-1}$ RG timescale beyond which the effective Hamiltonian $H_{0}(\lambda)$ approaches the prethermal fixed point. $t\ll t_{\text{th}}$ (defined below) Real timescale where the Floquet-Magnus expansion is a good approximation of the dynamics.
Prethermal regime: $\lambda_{\text{min}}\sim(1/J)\ln(\Omega/J)$ RG timescale where $\left\|{H_{1}(\lambda)}\right\|$} attains its minimal value after flowing near the prethermal fixed point. $t_{\text{eff}}(\lambda)\sim{\left\|{H_{1}(\lambda)}\right\|}^{-1}{\left\|{\hat{1}}\right\|}$ Real timescale upto which $H_{0}(\lambda)$ serves as a reliable approximation for dynamics.
Thermalization: $\lambda_{\text{th}}\gtrsim\lambda_{\text{min}}$ RG timescale where the system flows away from the prethermal fixed point. Thermalization: $t_{\text{th}}=t_{\text{eff}}(\lambda_{\text{min}})\sim(\Omega/J)^{\Omega/J}$ Real timescale beyond which the prethermal Hamiltonian $H_{0}(\lambda_{\text{min}})$ is no longer reliable.
Table 1: Glossary of flow renormalization group ($\lambda$) and real ($t$) time scales that describe the Floquet dynamics based on the Wegner-Wilson flow framework described in this work.